OpenStudy (anonymous):

multiply:

OpenStudy (anonymous):

\[2\sqrt{11} * 8\sqrt{22}\]

Parth (parthkohli):

I'll let you solve becca :D

OpenStudy (anonymous):

becca is answering questions now O_O

Parth (parthkohli):

Yeah I want her to reach 99 :D

OpenStudy (anonymous):

o_O

OpenStudy (anonymous):

i always give medals so i can help XD

Parth (parthkohli):

Okay becca doesnt know...I'll do it

Parth (parthkohli):

So, this'd be same as: \(\Large \color{Black}{\Rightarrow 2 \times \sqrt{11} \times 8 \times \sqrt{22} }\) \(\Large \color{Black}{\Rightarrow 16 \times \sqrt{11 \times 22} }\) \(\Large \color{Black}{\Rightarrow 16\sqrt{242} }\) Can you simplify the radical?

OpenStudy (anonymous):

\[8\sqrt{11}?\]

Parth (parthkohli):

No no

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 16\sqrt{242} = 16\sqrt{2 \times 11 \times 11} = 176\sqrt{2} }\)

OpenStudy (anonymous):

kk thanks